OpenStudy (anonymous):

While in the United States visiting her grandmother, Kendra copied the famous family apple pie recipe. When she returns to England, she plans on making the same apple pie for her friends. To do this, Kendra needs to convert the flour from cups to grams. She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams. Write a function that shows grams as a function of cups. Use g(x). I will give medal to whoever helps.

OpenStudy (anonymous):

At first I though it would be g(x)=128x but then I looked at it again and thought that maybe it was g(x)=128/x

OpenStudy (anonymous):

the right question is g(x) = 128x just try g(0.5) = 128*0.5 = 64 grams

OpenStudy (anonymous):

awesome can you help me with another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Evaluate g(x) for any value of x within the domain. Interpret the meaning of your solution in the context of this problem.

OpenStudy (anonymous):

part of the same question.

OpenStudy (anonymous):

The domain is \[0 \le x \le \infty\]

OpenStudy (anonymous):

so should that b emu answer or should I actually answer the function?

OpenStudy (anonymous):

*be my

OpenStudy (anonymous):

That's the answer, it's like, you have to interpret the meaning of g(x) in this context, which is calculate grams as a function of cups, what means transform cups into grams,

OpenStudy (anonymous):

Ok thanks you've helped me a lot!

OpenStudy (anonymous):

sorry man.. your domains is from - infinity to + infinity

OpenStudy (anonymous):

I misspelled up there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[-\infty \le x \le +\infty \]

OpenStudy (anonymous):

thats where your function can be true, in all this domain.. since you can also calculate negatives grams as negative cups

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

No problem